OpenStudy (anonymous):

If cos A = .5417 and A is an angle between 0 and 90, find A. Round to the nearest tenth of a degree

OpenStudy (binary3i):

A must be near pi/3

OpenStudy (anonymous):

i dont get you sorry can you explain

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

cos(pi/3)=1/2

OpenStudy (anonymous):

try use taylor

OpenStudy (anonymous):

i dont get the answer

OpenStudy (binary3i):

why dont you use calculator

OpenStudy (anonymous):

because my calculator is simple i cant type it in

OpenStudy (anonymous):

\[f(x)=f(x0)+(x-x0).f'(x0)+(x-x0)^{2}/2!.f''(x0)+....\]

OpenStudy (anonymous):

x0=pi/3;x=.5417

OpenStudy (anonymous):

would you mind telling me the answer

OpenStudy (anonymous):

i have no calculator now

OpenStudy (anonymous):

c:

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

0.5417=1/2+(x-pi/3).(-sen(pi/3))+(x-pi/3)²/2!.-(cos(pi/3))

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

your gona say im so dumb but i dnt know how

OpenStudy (anonymous):

its a second degree polynomial.Do you know baskara formula?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

A=57.20

OpenStudy (anonymous):

it's in degrees

OpenStudy (anonymous):

A. 32.0 B. 38.7 C. 51.3 but dude my answer choices are these D. 90.0